MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2016/12/09
OBSERVER: Alex Li
CONDITION: Windy & Cloudy
COMMENT: CCD temp reading was frozen before 11:30pm because the power supply output was off. The problem might have happened since 07/12 that's why no CCD temp plots of those days were uploaded.
Compressor gas pressure: 1.75 MPa
Expected Observation Time: START 23:03 END 04:16


Go to the Seeing Log

Go to the Observation Log

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB-000
LMC-000
SMC-000
TRANSIT-000
ALERT-000
GRB-000
TOTAL---000 (00h00m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
00000000000


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


TRANSIT FIELD STATISTICS

Field1Field2Field3
000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
23:03waitingWindy & Cloudy
00:00waitingWindy & Cloudy
01:00waitingWindy & Cloudy
02:00waitingWindy & cloudy
03:00waitingWindy & Cloudy
04:00waitingWindy & Cloudy
04:16waitingWindy & Cloudy


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day